In Desmos, type in the equation:

y=-3sqaure root of x+3


What is the domain of the graph in inequality notion?


1
Expert's answer
2021-03-18T08:07:30-0400

The function "y=\sqrt[3]{x+3}" is drawn in desmos as shown below



So, the domain of the graph in inequality notion is "-\infty \:<x<\infty"